# On the contraction properties of Sinkhorn semigroups

### v1

@@ 4.2 Norm contraction theorems
A proof of Theorem 4.2 is provided in Appendix C (on page C).
4.3 A Lyapunov stability theorem
Finally, we introduce a more tractable sufficient condition, in terms of the potential functions $(U,V,W)$. It enables us to obtain geometric rates for the convergence of the Sinkhorn iterations, as explicitly shown for the case of linear Gaussian models and finite mixture models in Sections 5.2 and 5.3, respectively. Let us assume that $W$ is locally bounded, $(U,V)\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y}))$ and $e^{-W}\in\mathbb{L}_{\infty}(\lambda\otimes\nu)$. In addition, assume that $U$ and $V$ have compact sub-levels sets with non empty interior and the measures $\lambda$ and $\nu$ have full support, i.e., they are finite on compact sets and strictly positive on non-empty open sets. Let us also denote
Condition (${\cal H}_{\delta}(U,V)$).
There is some $\delta\in]0,1[$ such that ${\cal U}_{\delta}:=\delta U-W^{V}\in{\cal B}_{\infty}(\mathbb{X})$ and ${\cal V}_{\delta}:=\delta V-W_{U}\in{\cal B}_{\infty}(\mathbb{Y})$. In addition, the functions $W_{U},W^{V}$ are locally bounded and
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee~{}\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty.$
@@ 4.3 A Lyapunov stability theorem
If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions $(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})$, where $U_{\star}=\inf_{\mathbb{X}}U(x)$ and $V_{\star}=\inf_{\mathbb{Y}}V(y)$. In addition, there exist some parameter $\rho\in[0,1[$ and some constant $c>0$ such that
@@
$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ ::=:= $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\cal B}_{\varphi}(\mathbb{X})$,}$
$\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ ::=:= $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\cal B}_{\psi}(\mathbb{Y})$},$ (43)
@@
$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2p+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right)$

### v2

@@ 4.4 A Lyapunov stability theorem
Condition 4.10 (${\cal H}^{\prime}_{\delta}(U,V)$).
There is some $\delta\in]0,1/2[$ such that ${\cal U}_{\delta}$ and ${\cal V}_{\delta}$ are locally bounded with compact sub-level sets and we have
$\lambda_{(1-2\delta)U}\left(1\right)\vee\nu_{(1-2\delta)V}\left(1\right)<\infty.$ (66)
We have
${\cal H}^{\prime}_{\delta}(U,V)\Longrightarrow{\cal H}_{\delta}(U,V).$ (67)
To verify this claim, note that the regularity condition on $({\cal U}_{\delta},{\cal V}_{\delta})$ ensures that
$(e^{{\cal U}_{\delta}},e^{{\cal V}_{\delta}})\in({\cal B}_{\infty}(\mathbb{X})\times{\cal B}_{\infty}(\mathbb{Y})).$
When ${\cal U}_{\delta}$ is locally bounded with compact sub-level sets $C_{{\cal U}}(r):=\{{\cal U}_{\delta}\leq r\}$ for any $0<\delta<1/2$ we have
$\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)$ $=$ $\lambda_{(1-2\delta)U}\left(e^{-{\cal U}_{\delta}}\right)$
$\leq$ $\left(e^{-r}+e^{c_{\delta}(r)}\right)\lambda_{(1-2\delta)U}\left(1\right)\quad\mbox{\rm with}\quad c_{\delta}(r):=\sup_{x\in C_{{\cal U}}(r)}|{\cal U}_{\delta}(x)|.$
In this context, we readily see that
[…]
This ends the proof of (67).
Remark 4.11.
Whenever $\|W\|<\infty$ we have $\|W_{U}\|\vee\|W^{V}\|<\infty$. In this case ${\cal H}_{\delta}(U,V)$ is satisfied as soon as there is some $\delta\in]0,1[$ such that
$\lambda_{(1-\delta)U}\left(1\right)\vee\nu_{(1-\delta)V}\left(1\right)<\infty.$ (69)
Remark 4.12.
Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions and we have $\delta U-W^{V}\geq{\cal U}_{\delta}^{-}$ and $\delta V-W_{U}\geq{\cal V}_{\delta}^{-}$, for some functions $({\cal U}_{\delta}^{-},{\cal V}_{\delta}^{-})$ with compact level sets. In this case, the functions $({\cal U}_{\delta},{\cal V}_{\delta})$ are locally bounded with compact level sets.
A Lyapunov stability theorem
@@
If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is also satisfied with $n_{0}=1$ for the Lyapunov functions
$(\varphi,\psi)=(e^{\delta(U-U_{\star})},e^{\delta(V-V_{\star})})\quad\mbox{with}\quad(U_{\star},V_{\star}):=(\inf_{\mathbb{X}}U,\inf_{\mathbb{Y}}V)\in(\mathbb{R}\times\mathbb{R}).$ (70)
In addition, there exist some parameter $\rho\in[0,1[$ and some constant $c>0$ such that
@@
The above theorem combined with the estimates (62) ensures that
$\sup_{n\geq n_{0}}\pi_{2n+1}(\varphi)$ $\leq$ $\epsilon~\nu_{V}(\psi)+c=\epsilon~e^{-\delta V_{\star}}~\nu_{(1-\delta)V}(1)+c$
$\sup_{n\geq n_{0}}\pi_{2n}(\psi)$ $\leq$ $\epsilon~\lambda_{U}(\varphi)+c=\epsilon~e^{-\delta U_{\star}}~\lambda_{(1-\delta)U}(1)+c$
with the parameters $(\epsilon,c)$ as in (60). The rescaling of the potential functions (70) is only made to have $\varphi\wedge\psi\geq 1$. Combining Theorem 4.13 with Theorem 4.8 we obtain the following corollary.
Corollary 4.14.
Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
[…]
with the functions $(\varphi,\psi)=(e^{\delta U},e^{\delta V})$.
The r.h.s. term in (72) can be estimate with (62). Also note that
[…]
This yields the exponential decays.
Corollary 4.15.
Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In this situation, there exists some parameters $n_{0}\geq 0$, $\rho\in[0,1[$ and $c_{1},c_{2}\geq 0$ such that for any $n\geq n_{0}$ we have
[…]
with the functions $(\psi,\varphi)$ as in Corollary 4.14.
4.5 Wasserstein exponential decays
Assume that the Lyapunov functions $(\psi,\varphi)$ defined in (70) satisfy the estimates
$e^{\delta U(x_{1})}+e^{\delta U(x_{2})}\geq\varrho_{\varphi}(x_{1},x_{2})\quad\mbox{\rm and}\quad e^{\delta V(y_{1})}+e^{\delta V(y_{2})}\geq\varrho_{\psi}(y_{1},y_{2})$ (74)
for some semi-metric $\varrho_{\varphi},\varrho_{\psi}$. In this case, by (10) there exists $c_{u},c_{v}>0$ such that
[…]
with the Wasserstein semi-metrics $\mathbb{W}_{\varrho_{\varphi}},\mathbb{W}_{\varrho_{\psi}}$ defined in (8). This yields the following Wasserstein estimates.
Corollary 4.16.
Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In addition, the Lyapunov functions $(e^{\delta U},e^{\delta V})$ satisfy (74) for some semi-metric $(\varrho_{\varphi},\varrho_{\psi})$. In this situation, there exists some $n_{0}\geq 0$ and $c_{u},c_{v}\geq 0$ such that for any $n\geq n_{0}$ we have
[…]
In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (72) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (70).
Assume there exists some semi-metric $\varrho_{U},\varrho_{V}$ and some constants $c_{u},c_{v}\in\mathbb{R}$ such that
$U(x_{1})+U(x_{2})$ $\geq$ $c_{u}+\varrho_{U}(x_{1},x_{2}),$
$V(y_{1})+V(y_{2})$ $\geq$ $c_{v}+\varrho_{V}(y_{1},y_{2}).$ (76)
In this case, by (35), for any $p\geq 1$ there exists some constant $c_{p}>0$ such that
[…]
with the $p$-Wasserstein semi-metrics $\mathbb{W}_{p,\varrho_{U}},\mathbb{W}_{p,\varrho_{V}}$ defined in (8). This yields the following Wasserstein estimates.
Corollary 4.17.
Assume ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$. In addition, the target potential functions $(U,V)$ satisfy (76) for some semi-metric $(\varrho_{U},\varrho_{V})$. In this situation, for any $p\geq 1$ there exists some $c_{p}\geq 0$ such that for any $n\geq n_{0}$ we have
[…]
In the above display, $n_{0}\geq 1$ and $\rho\in[0,1[$ are as in (72) and $(\psi,\varphi)$ are the Lyapunov functions that depend on $\delta$ defined in (70).
@@ 4.6 Entropy exponential decays
$\lim_{n\rightarrow\infty}U_{2n}=\mathbb{U}$ $:=$ $U_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p+1}}{d\lambda_{U}}}\quad\mbox{\rm in ${\cal B}_{\varphi}(\mathbb{X})$,}$
$\lim_{n\rightarrow\infty}V_{2n}=\mathbb{V}$ $:=$ $V_{0}+\sum_{p\geq 0}\log{\frac{d\pi_{2p}}{d\nu_{V}}}\quad\mbox{\rm in ${\cal B}_{\psi}(\mathbb{Y})$},$ (78)
@@
$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2p+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right).$